A mixture consisting of 5.4 g CO and 9.2 g SO2, two atmospheric pollutants, has a pressure of 0.300 atm when placed in a sealed container. What is the partial pressure of the SO2?
1
Expert's answer
2019-10-14T14:31:39-0400

According to Dalton's law, partial pressure of the SO2 is equal to

"P_{SO_2}=x_{SO_2} \cdot P_{total}"

where xSO2 is mole fraction of SO2 and Ptotal is total pressure of gases.

As total pressure is known, we should find mole fraction of SO2.

"x_{SO_2}=\frac{n(SO_2)}{n(SO_2)+n(CO)}=\frac{\frac{m(SO_2)}{M(SO_2)}}{\frac{m(SO_2)}{M(SO_2)}+\frac{m(CO)}{M(CO)}}=\frac{\frac{9.2g}{64 \frac{g}{mol}}}{\frac{9.2g}{64 \frac{g}{mol}}+\frac{5.4g}{28 \frac{g}{mol}}}=0.427"

Thus, partial pressure of the SO2 is equal to

"P_{SO_2}=0.427 \cdot 0.300 atm = 0.128 atm"